# Descrip3D: Enhancing Large Language Model-based 3D Scene Understanding with Object-Level Text Descriptions

### v1

@@ Appendix C Additional Quantitative Results
We evaluate our method using the standard metrics established in the original papers for each 3D scene-language dataset. To thoroughly assess the effectiveness of our approach, we perform extensive comparisons against a diverse set of baselines across multiple benchmarks. To complement the main results, we report additional evaluation metrics on the same datasets (ScanRefer, Multi3DRefer, and ScanQA) used in the main paper. The results, summarized in Table 8 (ScanRefer), Table 9 (Multi3DRefer), and Table 10 (ScanQA), show our method consistently outperforms prior approaches across grounding and question answering tasks. On ScanRefer, Descrip3D achieves the highest overall accuracy. On Multi3DRefer, it leads in almost all grounding settings, with the best overall F1 scores. On ScanQA, it outperforms baselines in nearly all language metrics, including ROUGE-L, METEOR, and CIDEr. These results confirm the effectiveness of incorporating object-level textual descriptions through dual-level integration for 3D vision-language tasks.
Appendix D Prompt Template
@@
User: What is the ID of the object that matches the description "this is a long table. it is surrounded by chairs"?
Assistant: <OBJ023>
@@
ScanRefer [4] ECCV20 76.33 53.51 32.73 21.11 41.19 27.40
TGNN [24] AAAI21 68.61 56.80 29.84 23.18 37.37 29.70
SAT [51] ICCV21 73.21 50.83 37.64 25.16 44.54 30.14
InstanceRefer [55] ICCV21 75.72 64.66 29.41 22.99 38.40 31.08
3DVG-Transformer [59] ICCV21 81.93 60.64 39.30 28.42 47.57 34.67
MVT [25] CVPR22 77.67 66.45 31.92 25.26 40.80 33.26
3D-SPS [32] CVPR22 84.12 66.72 40.32 29.82 48.82 36.98
ViL3DRel [7] NeurIPS22 81.58 68.62 40.30 30.71 47.94 37.73
3DJCG [3] CVPR22 83.47 64.34 41.39 30.82 49.56 37.33
D3Net [5] ECCV22 72.04 30.05 37.87
BUTD-DETR [26] ECCV22 84.2 66.3 46.6 35.1 52.2 39.8
HAM [6] ArXiv22 79.24 67.86 41.46 34.03 48.79 40.60
3DRP-Net [43] EMNLP23 83.13 67.74 42.14 31.95 50.10 38.90
3D-VLP [27] CVPR23 84.23 64.61 43.51 33.41 51.41 39.46
EDA [47] CVPR23 85.76 68.57 49.13 37.64 54.59 42.26
M3DRef-CLIP [58] ICCV23 85.3 77.2 43.8 36.8 51.9 44.7
3D-VisTA [62] ICCV23 81.6 75.1 43.7 39.1 50.6 45.8
ConcreteNet [41] ECCV24 86.40 82.05 42.41 38.39 50.61 46.53
DORa [46] ArXiv24 52.80 44.80
Chat-Scene [22] NeurIPS24 89.59 82.49 47.78 42.90 55.52 50.23
@@
Table 8: Performance comparison on the validation set of ScanRefer [4].
@@
3DVG-Trans+ [59] ICCV21 87.1 45.8 16.7 26.5 25.5
D3Net (Grounding) [5] ECCV22 81.6 32.5 23.3 35.0 32.2
3DJCG (Grounding) [3] CVPR22 94.1 66.9 16.7 26.2 26.6
M3DRef-CLIP [58] ICCV23 81.8 39.4 53.5 47.8 34.6 30.6 43.6 37.9 42.8 38.4
Chat-Scene [22] NeurIPS24 90.3 62.6 82.9 75.9 49.1 44.5 45.7 41.1 57.1 52.4
@@
Table 9: Performance comparison on the validation set of Multi3DRefer [58].
@@
ScanQA [2] CVPR22 21.05 30.24 20.40 15.11 10.08 33.33 13.14 64.86
3D-VLP [27] CVPR22 21.65 30.53 21.33 16.67 11.15 34.51 13.53 66.97
3D-LLM [20] NeurIPS23 20.5 39.3 25.2 18.4 12.0 35.7 14.5 69.4
LL3DA [8] CVPR24 13.53 37.31 15.88 76.79
LEO [23] ICML24 11.5 39.3 16.2 80.0
Scene-LLM [14] WACV25 27.2 43.6 26.8 19.1 12.0 40.0 16.6 80.0
Chat-Scene [22] NeurIPS24 21.62 43.20 29.06 20.57 14.31 41.56 18.00 87.70
@@
Table 10: Performance comparison on the validation set of ScanQA [2].

### v2

@@ Appendix F Additional Quantitative Results
We evaluate our method using the standard metrics established in the original papers for each 3D scene-language dataset. To thoroughly assess the effectiveness of our approach, we perform extensive comparisons against a diverse set of baselines across multiple benchmarks. To complement the main results, we report additional evaluation metrics on the same datasets (ScanRefer, Multi3DRefer, and ScanQA) used in the main paper. The results, summarized in table˜12 (ScanRefer), table˜13 (Multi3DRefer), and table˜14 (ScanQA), show our method consistently outperforms prior approaches across grounding and question answering tasks. On ScanRefer, Descrip3D achieves the highest overall accuracy. On Multi3DRefer, it leads in almost all grounding settings, with the best overall F1 scores. On ScanQA, it outperforms baselines in nearly all language metrics, including ROUGE-L, METEOR, and CIDEr. These results confirm the effectiveness of incorporating object-level textual descriptions through dual-level integration for 3D vision-language tasks.
Appendix G Prompt Template
@@
Figure 8: IoU distribution comparison between Chat-Scene and Descrip3D on (a) ScanRefer and (b) Multi3DRefer. The histograms show the per-sample IoU distributions, with dashed vertical lines indicating the 0.25 and 0.5 thresholds commonly used for visual grounding. Compared to Chat-Scene, Descrip3D consistently reduces the concentration of low-IoU cases and increases the density of high-IoU predictions, demonstrating more robust performance across the full distribution.
@@
User: What is the ID of the object that matches the description "this is a brown chair. It is to the right of another chair near the end of the table."? [Generated description (may be noisy)] <OBJ002>: there is a <OBJ002> positioned next to a <OBJ003>…
Assistant: <OBJ002>
@@
ScanRefer [chen2020scanrefer] ECCV20 76.33 53.51 32.73 21.11 41.19 27.40
TGNN [huang2021text] AAAI21 68.61 56.80 29.84 23.18 37.37 29.70
SAT [yang2021sat] ICCV21 73.21 50.83 37.64 25.16 44.54 30.14
InstanceRefer [yuan2021instancerefer] ICCV21 75.72 64.66 29.41 22.99 38.40 31.08
3DVG-Transformer [zhao20213dvg] ICCV21 81.93 60.64 39.30 28.42 47.57 34.67
MVT [huang2022multi] CVPR22 77.67 66.45 31.92 25.26 40.80 33.26
3D-SPS [luo20223d] CVPR22 84.12 66.72 40.32 29.82 48.82 36.98
ViL3DRel [chen2022language] NeurIPS22 81.58 68.62 40.30 30.71 47.94 37.73
3DJCG [cai20223djcg] CVPR22 83.47 64.34 41.39 30.82 49.56 37.33
D3Net [chen2022d3net] ECCV22 72.04 30.05 37.87
BUTD-DETR [jain2022bottom] ECCV22 84.2 66.3 46.6 35.1 52.2 39.8
HAM [chen2022ham] ArXiv22 79.24 67.86 41.46 34.03 48.79 40.60
3DRP-Net [wang20233drp] EMNLP23 83.13 67.74 42.14 31.95 50.10 38.90
3D-VLP [jin2023context] CVPR23 84.23 64.61 43.51 33.41 51.41 39.46
EDA [wu2023eda] CVPR23 85.76 68.57 49.13 37.64 54.59 42.26
M3DRef-CLIP [zhang2023multi3drefer] ICCV23 85.3 77.2 43.8 36.8 51.9 44.7
3D-VisTA [zhu20233d] ICCV23 81.6 75.1 43.7 39.1 50.6 45.8
ConcreteNet [unal2023three] ECCV24 86.40 82.05 42.41 38.39 50.61 46.53
DORa [wu2024dora] ArXiv24 52.80 44.80
Chat-Scene [huang2024chat] NeurIPS24 89.59 82.49 47.78 42.90 55.52 50.23
@@
Table 12: Performance comparison on the validation set of ScanRefer [chen2020scanrefer].
@@
3DVG-Trans+ [zhao20213dvg] ICCV21 87.1 45.8 16.7 26.5 25.5
D3Net (Grounding) [chen2022d3net] ECCV22 81.6 32.5 23.3 35.0 32.2
3DJCG (Grounding) [cai20223djcg] CVPR22 94.1 66.9 16.7 26.2 26.6
M3DRef-CLIP [zhang2023multi3drefer] ICCV23 81.8 39.4 53.5 47.8 34.6 30.6 43.6 37.9 42.8 38.4
Chat-Scene [huang2024chat] NeurIPS24 90.3 62.6 82.9 75.9 49.1 44.5 45.7 41.1 57.1 52.4
@@
Table 13: Performance comparison on the validation set of Multi3DRefer [zhang2023multi3drefer].
@@
ScanQA [azuma2022scanqa] CVPR22 21.05 30.24 20.40 15.11 10.08 33.33 13.14 64.86
3D-VLP [jin2023context] CVPR22 21.65 30.53 21.33 16.67 11.15 34.51 13.53 66.97
3D-LLM [hong20233d] NeurIPS23 20.5 39.3 25.2 18.4 12.0 35.7 14.5 69.4
LL3DA [chen2024ll3da] CVPR24 13.53 37.31 15.88 76.79
LEO [huang2023embodied] ICML24 11.5 39.3 16.2 80.0
Scene-LLM [fu2024scene] WACV25 27.2 43.6 26.8 19.1 12.0 40.0 16.6 80.0
Chat-Scene [huang2024chat] NeurIPS24 21.62 43.20 29.06 20.57 14.31 41.56 18.00 87.70
@@
Table 14: Performance comparison on the validation set of ScanQA [azuma2022scanqa].
Appendix H IoU Variance and Distribution
To provide a more fine-grained view of grounding robustness beyond thresholded metrics (Acc@0.25, Acc@0.50), we report IoU distributions and variance. fig.˜8 shows per-sample IoU histograms for ScanRefer and Multi3DRefer, both following the typical bimodal pattern with concentrations near IoU = 0 (failure) and IoU = 1 (successful grounding). Compared to Chat-Scene, Descrip3D consistently reduces the density of near-zero IoUs and increases the high-IoU mass. The measured variance is stable across datasets: ScanRefer variance $\approx$ 0.19 (std $\approx$ 0.44) and Multi3DRefer variance $\approx$ 0.20 (std $\approx$ 0.44). These results demonstrate that improvements are not driven by a handful of outlier cases but reflect consistent robustness across samples.